OpenStudy (anonymous):

I need some help understanding how to solve this equation: (-2x^2y-4)(4x^-1y^3)^3

OpenStudy (anonymous):

whenever u have negative exponents u have to rewrite the number as a fraction. for example: 3^ -2 = 1 / 3^2 = 1/9

OpenStudy (anonymous):

Thanks:)

OpenStudy (anonymous):

well do u understand how to do it now?

OpenStudy (anonymous):

I do thanks :)

OpenStudy (anonymous):

good. u wanna give me a medal? :)